Hi, how can I solve this equation using elimination?
4x + 5y = 57
7x - 2y = 46

Answer by checkley79(3341)

4x + 5y = 57 multiply by 2.
7x - 2y = 46 multiply by 5 & add.
8x+10y=114
35x-10y=230
--------------------
43x=344
x=344/43
x=8 ans.
4*8+5y=57
32+5y=57
y=57-32
y=25/5
y=5 ans.
proof:
7*8-2*5=46
56-10=46
46=46